1. Find all solutions to the equation 6sin(x-1)=2 for 0 < < 2π Give your answers to 3 decimal places.
Answer #1

For any further assistance feel free to ask in comment.6sin (n-1)=2 sin (n-1)= 2 , OC Xc2a sin(x-1) = 1 sin (x-1) = sin c General values AL ña +619 10.3390) = x= lina tel (0.3398)